OpenStudy (anonymous):

Would someone be willing to check my answers please? I promise to fan and medal!

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

1. A playground is being designed where children can interact with their friends in certain combinations. If there is 1 child, there can be 0 interactions. If there are 2 children, there can be only 1 interaction. If there are 3 children, there can be 4 interactions. If there are 4 children, there can be 9 interactions. How many interactions will there be for 8 children? (1 point) A. 49*** B. 64 C. 36 D. 40

OpenStudy (anonymous):

i will check

OpenStudy (anonymous):

you are right

OpenStudy (anonymous):



OpenStudy (anonymous):

its a check mark

OpenStudy (anonymous):

2. A playground is being designed where children can interact with their friends in certain combinations. If there is 1 child, there can be 0 interactions. If there are 2 children, there can be only 1 interaction. If there are 3 children, there can be 4 interactions. If there are 4 children, there can be 9 interactions. Which recursive equation represents the pattern? (1 point) an = an – 1 + 3n – 2 an = an – 1 + 2n – 3*** an = an – 1 + (3n – 2) an = an – 1 + (2n – 3)

OpenStudy (anonymous):

i have done these before

OpenStudy (anonymous):

oh really? cool!

OpenStudy (anonymous):

and yes thats right aggain

OpenStudy (anonymous):

sweet! Sheila purchased a plot of land for $23,500. The land appreciates about 1.9% each year. What is the value of the land after seven years? (1 point) $20,547.12 $23,761.70 $26,625.50*** $26,809.40

OpenStudy (anonymous):

whaat grade are you in

OpenStudy (anonymous):

I'm in 11th

OpenStudy (anonymous):

ok and i have answer key for this and thats right and my brother asked me to tell you to try looking at google

OpenStudy (anonymous):

google.com

OpenStudy (anonymous):

lol i did. I was just making sure...

OpenStudy (anonymous):

brb my bro will help with more if you need it dont worry hes nice

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

will you help with my math plz

OpenStudy (anonymous):

I'm not very good at math...

OpenStudy (anonymous):

What is the approximate area of the figure? A. 94.6 square feet B. 114.3 square feet C. 84.8 square feet D. 49.6 square feet

OpenStudy (anonymous):

thats the figure link

OpenStudy (anonymous):

I'm not sure...

OpenStudy (anonymous):

Yes that is correct. That is a simple interest problem. Simple Interest = Principal amount * Interest Rate * period of time = 23500 * 1.9% * 7 = 23500 * 0.019 * 7 = 3125.5 So the internet she earns on the land after 7 years is 3125.5. Therefore, the total worth of the land after 7 years is 23500 + 3125.5 = 26625.5

OpenStudy (anonymous):

Sorry I meant "interest" not "internet" in the above answer lol.

OpenStudy (anonymous):

genius 12 you got that from google i just saw that

OpenStudy (anonymous):

0,o? I beg your pardon? I just solved that right now.

OpenStudy (anonymous):

Those are my own words and I had no idea that this problem is on Google? 0.0

OpenStudy (anonymous):

nope i just saw it sorry

OpenStudy (anonymous):

not your own work

OpenStudy (anonymous):

sorry about that that was my brother

OpenStudy (anonymous):

hey foreverbrianna1 i thout you said a promisd fan and medal i helped with all qs

OpenStudy (anonymous):

done and done thank you so much! I really appreciate all your help :)

OpenStudy (anonymous):

thank you